The orientation of small satellites is often controlled using torque from current-carrying coils in Earth's magnetic field. Suppose a multiturn coil has a cross-sectional area of 6.96 ✕ 10−4 m2, dissipates 0.300 W of electrical power from a 5.50 V power supply, and provides a magnetic moment of magnitude 0.0400 A · m2.

(a) Find the coil current I (in A).

A

(b) Calculate the number of turns in the coil.

turns

(c) Calculate the maximum magnitude of torque (in N · m) if Earth's magnetic field has magnitude 3.60 ✕ 10−5 T at the satellite's location.

N · m
Answer #1

(a) I = P/V= (0.300 W)/(5.50 V)

= 0.0545 A

(b) M = NIA

.0400 =N× 0.0545 ×(6.96×10-4)

N = 1054 turns

(c) T= MB sine

Tmar = MB = (0.0400)×(3.60×10-5)

= 1.44×10-6 N.m